if w is subspace of a vector space V over field F such that dim v=5 and dim w=2 then dim A(w) =



1
Expert's answer
2021-08-04T13:11:59-0400

"\text{Since w is a subspace of v, we have that,}\\dim(w)+dim(A(w))=dim(v)
\\\text{$dim(w) = 5-2=3$}"